# Trading Carbon for Physics: On the Resource Efficiency of Machine Learning for Spatio-Temporal Forecasting

### v1

@@ 5 Results
5.1 Harmonic Oscillator Toy Experiment
The results in Fig. 2 show that models with non-periodic activations perform poorly outside the training domain (extrapolation). Periodic activations (Sine, Snake) offer modest improvements, with the model using Snake activation function performing best. The PINN achieved superior results in both interpolation and extrapolation tasks; most notably for the damped harmonic oscillator shown in Fig 5. It remained stable and accurate even as prediction range increased and dynamics grew more complex.
5.2 Viscous Burgers’ Equation Experiment
For the viscous Burgers’ equation experiment, the comparison with NN, U-PINN, and S-PINN show interesting trends. Both PINNs achieve consistently lower test losses at lower carbon footprints compared to the supervised NN (see Fig. 1-b). Although the NN has the lowest carbon footprint for fixed epochs and data size, achieving a given accuracy is up to $10$ times more carbon-efficient with the PINNs. This S-PINN, which mixes data and physics, dominates the Pareto front in Fig. 1-b, suggesting that combining physics and data gives the model which balances efficacy and efficiency.
5.3 Incompressible Shear Flow Experiment
The first two simple experiments discussed above establish key insights into how physics can affect the model performance and resource consumption for specific tasks. In the incompressible shear flow experiment, we extend this analysis to a more complex task using a wide spectrum of models.
@@ 5.3 Incompressible Shear Flow Experiment
We first assess predictive performance using Pearson correlation $r$ at each step of a 20-step rollout, see Fig. 6. All eight models (see Table 1) achieve high accuracy for the first few steps (r $\approx$ 1), after which performance declines at different rates. Over longer horizons, UNet-CN, UNet-CN-CP, and UFNO emerge as the strongest performers.
Figure 6: Pearson r over 20 rollout steps. Values are averaged over the four fields: velocity components, tracer, and pressure. Models labels are ordered from no bias (top) to strong bias (bottom).
Figure 7 illustrates these stability differences visually. UNet and UNet-CP deteriorate rapidly, accumulating distortions within just a few steps. The modernised variants (UNet-CN, UNet-CN-CP) show marked improvements: UNet-CN captures fine-scale detail but gradually disrupts large-scale dynamics, while UNet-CN-CP better preserves both large and fine-scale structure at the cost of high-frequency artifacts. FM exhibits good temporal stability, though it tends to exaggerate some fine-scale structures. FNO and TFNO maintain the dominant large-scale dynamics but exhibit a systematic drift toward higher values, fail to capture fine-scale variability and eventually introduce distortions. UFNO better recovers both large- and small-scale structures, though these gradually smooth out over time.
Figure 7: Example rollout predictions for one of the four fields (tracer). Target trajectory (top row) compared with model predictions shown every third step of a 20-step rollout, illustrating differences in long-horizon stability and fidelity. Models labels are ordered from no bias (top) to strong bias (bottom). Results for the three other fields are provided in Appendix G.
Training Cost.
@@
Inference Cost.
To provide a more complete evaluation, we also account for inference costs of all the trained models. This can be useful in choosing spatio-temporal forecasting models as inference costs can dominate over training or development costs.
Figure 8: Predictive performance and carbon footprint. a: Predictive accuracy (Pearson r) versus training emissions (kgCO2eq); each point represents every second step of a 20-step rollout (lighter shades indicate earlier steps), averaged over four fields. b: Training emissions (kgCO2eq) from a single run per model. Models are ordered from no bias (left) to strong bias (right). c: One-step inference emissions (gCO2eq) from three repeated runs over the full test set; error bars indicate variability across repetitions. Note that inference costs also include the initial model loading cost.
[…]
Padding choices further illustrate the sensitivity of emissions to architectural design. Switching from zero to circular padding increases training emissions for both UNet and UNet-CN. At inference, the effects diverge: circular padding slightly increases the footprint for UNet, while significantly reducing it for UNet-CN. This highlights how even minor inductive design choices – in this case enforcing circular padding to reflect periodic boundaries – can shift the efficacy–efficiency balance.
Table 6 in App. G reports both training and inference footprints alongside run times. While runtime and energy consumption/ carbon emissions are closely linked, they are not strictly proportional, as some operations could take longer but use less energy, and vice-versa (Henderson et al., 2022; Bakhtiarifard et al., 2024). Additional results for these models can be found in App. G.
@@ 6 Discussions
Our study demonstrates that physics inductive biases impact the trade-off between efficacy and efficiency. In simple settings, such as the harmonic oscillator, adding periodic structure or a physics-informed loss enables extrapolation beyond the training range. With viscous Burgers’ equation, combining data with physics yields the model that dominates the Pareto front, achieving lower test errors at lower carbon cost compared to the purely data-driven baseline.
@@
In the incompressible shear flow setting, we observe clear differences in how the physics priors affect both predictive performance and carbon footprint across training and inference. Models with strong physics inductive biases achieve the lowest training footprints, but this comes at the expense of higher inference costs compared to U-net variants. UFNO achieves one of the most favourable overall profiles, combining strong predictive accuracy with a low training footprint. Its inference cost, however, remains nearly twice that of the best-performing U-Net variant (UNet-CN-CP). The results underscore the importance of evaluating training and inference together. A model that trains cheaply may prove inefficient when repeatedly deployed, while others entail higher upfront cost but amortise more effectively in deployment. Amortising development costs across the model lifecycle is one way to factor in these considerations (Cottier et al., 2024).
FM stands out as an outlier: while competitive in predictive performance, its reliance on numerical integration makes inference substantially more expensive. This is in spite of it operating on downsampled data, as the original resolution was computationally prohibitive (Dao et al., 2023).
@@
Several key points follow from these results. First, model selection cannot be based on efficacy alone, as models with similar predictive accuracy can differ substantially in carbon emissions. Second, once both training and inference costs are accounted for, identifying a single “best” model becomes less straightforward, as illustrated by the incompressible shear flow experiment. A model with a small carbon footprint during training may be unsuitable in deployment, making carbon-aware evaluation across the full lifecycle essential.
@@
The implications extend well beyond our test-bed. Many scientific and engineering problems are governed by well-understood structures, and overlooking them is wasteful. At the same time, tailoring models too closely to the specific physics can limit applicability and can be challenging for complex systems. This motivates the development of general-purpose models such as FNOs and FM, which allow physics priors to be incorporated more easily while retaining flexibility across tasks. However, as demonstrated by FM, alignment with the underlying physics does not guarantee carbon efficiency. Physics should not be incorporated at any cost, but only where it provides clear benefits in terms of predictive performance or carbon efficiency. While our discussion centres on physics-inductive biases, the same logic extends to other forms of prior knowledge: any regularising bias that constrains the learning task appropriately can potentially offer similar trade-offs between efficacy and efficiency.
@@
Ultimately, our results argue for a shift in mindset. Progress in ML should not be equated with ever-larger datasets, models, and compute budgets, but rather evaluated in terms of the value gained relative to the cost incurred. Learning everything from data while ignoring domain knowledge reflects an abundance mindset, in which resources are treated as unlimited. A more sustainable alternative is to exploit what we already know by embedding this knowledge into models. Trading carbon for physics offers a principled path towards architectures that leverage prior knowledge, reduce reliance on brute-force compute, and achieve more favourable efficacy-efficiency trade-offs.
@@
This study has some limitations. First, all models were trained without extensive hyperparameter tuning. This choice emphasises architectural comparison over peak performance, trading statistical robustness for a broader perspective on performance–carbon trade-offs. In practice, however, hyperparameter search can dominate computational cost, and complex models typically require many more training runs during development than simpler baselines. Second, we restricted our study to spatio-temporal forecasting. While this limits the scope of the contribution, the central message about taking efficiency considerations into account by including appropriate inductive biases is applicable across ML tasks. Finally, our measurements do not account for hardware specific optimisations, where certain operations are more efficiently executed on specific accelerators. This, for instance, could yield an advantage to deep learning models compared to PDE solvers.
@@ 7 Conclusions
The primary message in this work, backed by empirical evidence, is to increase resource-awareness when designing ML experiments. While this can be accomplished in several ways (Bartoldson et al., 2023), we focus on the role of physics inductive biases and illustrate the different decisions practitioners have to consider. Through multiple experiments from simple oscillators to complex shear flow, we showed how physics-inductive biases affect the trade-off between predictive performance and carbon footprint.
This work advances a perspective grounded in resource-awareness and calls for a shift in mindset. As a community, we should discourage the pursuit of marginal improvements in performance that come from disproportionality higher resource costs. With our work we have made the case to encourage resource-awareness by utilising existing domain knowledge and/or other inductive biases.
Acknowledgements
@@
Appendix A Flow Matching
We condition on two consecutive states $(u_{0},u_{1}):=(u_{k},u_{k+1})$ such that the path interpolates between current state $u_{0}$ and the target state $u_{1}$. Following Lim et al. (2025), we adopt their interpolation path:
$p_{\tau}(u\mid u_{0},u_{1})=\mathcal{N}\!\left(u\;\middle|\;(1-\tau)u_{0}+\tau u_{1},\;\big(\sigma_{\min}^{2}+\sigma^{2}\tau(1-\tau)\big)I\right),$ (2)
with $\sigma_{\min}=10^{-3}$ and $\sigma=10^{-1}$, as recommended for Navier–Stokes data. The mean interpolates linearly between $u_{0}$ and $u_{1}$, while the variance follows a Brownian-bridge profile (maximal halfway at $\tau=0.5$, minimal at endpoints at $\tau=0$ and $\tau=1$).
The corresponding vector field is:
$v_{\tau}(u\mid u_{0},u_{1})=u_{1}-u_{0}+\frac{\sigma^{2}}{2}\,\frac{1-2\tau}{\sigma_{\min}^{2}+\sigma^{2}\tau(1-\tau)}\Big(u-\big((1-\tau)u_{0}+\tau u_{1}\big)\Big).$ (3)
The parametric vector field $v_{\tau}^{\theta}$ is trained to approximate $v_{\tau}$ by minimizing:
$\mathcal{L}(\theta)=\mathbb{E}_{\tau,\;p_{\tau}(u\mid u_{0},u_{1}),\;q(u_{0},u_{1})}\Big\|v_{\tau}^{\theta}(u)-v_{\tau}(u\mid u_{0},u_{1})\Big\|^{2}.$ (4)
Here $\tau\sim\mathcal{U}[0,1]$, $u\sim p_{\tau}(u\mid u_{0},u_{1})$, and $(u_{0},u_{1})\sim q(u_{0},u_{1})$, where $q$ denotes the empirical distribution of consecutive state pairs in the dataset.
In addition to the conditioning pair $(u_{0},u_{1})$, we provide temporal context by sampling one additional state $u_{c}\in\{u_{k-3},u_{k-2},u_{k-1}\}$, giving a maximum horizon $h=4$ consistent with FNOs and U-nets. All inputs are downsampled by a factor of 16 via average pooling, and predictions are upsampled to the original resolution with bilinear interpolation.
During inference, given a state $u_{k}$, the learned flow is simulated by numerically integrating $v_{\tau}^{\theta}$ from $\tau=0$ to $\tau=1$. We use Euler’s method with $10$ integration steps, which produces intermediate states approximating the continuous trajectory. The final iterate is then upsampled to the original resolution, yielding the forecast $\hat{u}_{k+1}$.
Appendix B Toy Experiment on Harmonic Oscillator
We design a toy experiment to study the effect of physics inductive biases for extrapolation in dynamical systems. The task is to model the trajectory of a harmonic oscillator, comparing standard feedforward networks with variants that incorporate different inductive biases. Specifically, we evaluate four networks with identical architecture but different activation functions, ReLU, Tanh, Sine, and Snake (Ziyin et al., 2020), where:
$\text{Sine}(x)=\sin(x),\qquad\text{Snake}(x)=x+\sin^{2}(x).$ (5)
The Sine and Snake activations introduce oscillatory biases aligned with the dynamics of the oscillator. In addition, we include a PINN that augments the data loss with a PDE residual while using Snake activations.
The governing dynamics are given by the damped harmonic oscillator:
$m\frac{d^{2}u}{dt^{2}}+\mu\frac{du}{dt}+ku=0\penalty 10000\ ,$ (6)
$u(0)=1,\penalty 10000\ \penalty 10000\ \mathrm{and}\penalty 10000\ \penalty 10000\ \frac{du}{dt}=0.$ (7)
The damping ratio is defined as $\delta=\mu/(2m)$. We adopt non-dimensionalized units with $m=1$ and $\omega_{0}=\sqrt{k/m}=1$. We consider both the undamped case ($\delta=0$) and a damped case ($\delta=0.02$).
Training data are sampled from $t\in[-5\pi,5\pi]$ with Gaussian noise ($\mu=0$$\sigma=0.3$). The purely supervised networks use $10,000$ supervised samples, while the PINN uses $100$ points ($1\%$) for the physics loss. Test data are noise-free and sampled over $t\in[5\pi,25\pi]$ to assess extrapolation. The points for evaluating the physics-based loss are sampled across the entire domain.
The purely supervised networks minimize the MSE:
$\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2},$ (8)
where $u$ denotes the ground truth and $\hat{u}$ the model predictions.
The PINN combines the same data term with a differential equation (DE) residual:
$\mathcal{L}_{\text{PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}+\frac{1}{N_{\text{de}}}\sum_{j=1}^{N_{\text{de}}}\left(\frac{d^{2}\hat{u}_{j}}{dt^{2}}+\mu\frac{d\hat{u}_{j}}{dt}+k\hat{u}_{j}\right)^{2}.$ (9)
Further architectural and optimisation details are given in Sec. 4.1.
Appendix C Loss Functions Used in the viscous Burgers’ Equation Experiment
The loss function for the purely data-driven NN, denoted $\mathcal{L}_{\text{NN}}$, is the MSE loss given by:
$\mathcal{L}_{\text{NN}}$ $=\mathcal{L}_{\text{data}}=\dfrac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}.$ (10)
Here $u$ and $\hat{u}$ denote the ground truth and the predictions.
The U-PINN optimizes solely using the governing equations including those that describe the boundary and initial conditions to ensure a unique solution. Its loss function, $\mathcal{L}_{\text{U-PINN}}$, is composed of the residuals of the PDE and the deviations from the initial and boundary conditions, expressed as:
@@ Appendix C Loss Functions Used in the viscous Burgers’ Equation Experiment
$=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left(\frac{\partial\hat{u}_{i}}{\partial t}+\hat{u}_{i}\frac{\partial\hat{u}_{i}}{\partial x}-\nu\frac{\partial^{2}\hat{u}_{i}}{\partial x^{2}}\right)^{2}$ (12)
@@
The S-PINN combines both data and physics losses. Its total loss function, $\mathcal{L}_{\text{S-PINN}}$ is given by:
$\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{NN}}+\mathcal{L}_{\text{U-PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}.$ (13)
Appendix D Incompressible Shear Flow Dataset
This experiment is based on data from The Well (Ohana et al., 2024), consisting of 2D incompressible Navier–Stokes dynamics with an additional passive tracer:

### v2

@@ 4 Results
Figure 4: Pearson r over 20 rollout steps. Values are averaged over trajectories and fields. Models labels are ordered from no bias (top) to strong bias (bottom).
@@
Carbon Footprint.
We next evaluate the carbon footprint and runtime associated with both training and inference (Table 2). Considering carbon efficiency (CO2eq per epoch), FNO and TFNO are roughly as carbon-intensive as the baseline UNet, while the other U-net variants are considerably more intensive. Within the U-net family, both the ConvNext design and circular padding make the model more carbon-intensive, with the exception of UNet-CN-CP, which is more efficient relative to UNet-CN. This indicates that architectural design choices can interact with computational efficiency in non-trivial ways. UFNO stands out as the by far most carbon-efficient model.
Table 2: Carbon footprint during training and inference. Reported CO2eq emissions and runtime for all models, measured with CarbonTracker. Inference results correspond to one-step forecasts over the full test set (4,704 samples) and include the initial model loading cost. Each inference experiment was repeated three times; CO2eq is reported as mean and standard deviation, and runtimes as mean values.
Training Inference (one-step)
Model CO2eq (kg) Time (h) # Epochs CO2eq/Epoch (g) CO2eq (g) Time (min)
UNet $2.72$ $12.7$ $43$ $63$ $3.68\pm 0.03$ $1.0$
UNet-CN $4.71$ $15.9$ $27$ $174$ $7.77\pm 0.05$ $1.9$
UNet-CP $3.81$ $14.2$ $35$ $109$ $4.06\pm 0.05$ $1.0$
UNet-CN-CP $5.89$ $20.9$ $45$ $131$ $4.87\pm 0.02$ $1.2$
FNO $0.80$ $4.0$ $12$ $67$ $7.87\pm 0.94$ $2.0$
TFNO $1.09$ $5.3$ $16$ $68$ $7.21\pm 0.12$ $1.8$
UFNO $1.12$ $3.6$ $30$ $37$ $9.00\pm 0.88$ $2.4$
Figure 5: a: Training emissions (kgCO2eq) from a single run per model. b: One-step inference emissions (gCO2eq) from three repeated runs over the full test set; error bars indicate variability across repetitions. Note that inference costs also include the initial model loading cost. c: Carbon footprint of inference during rollouts. Reported CO2eq emissions for 24 autoregressive rollouts of 50 steps. Inference costs include the initial model loading cost. Models are ordered from no bias (left) to strong bias (right).
Fig. 5 summarises emissions for training (a), one-step inference (b), and rollout inference (c). Comparing total training emissions, the carbon footprints of the FNO variants are between 2.4 and 7.4 times lower than those of the U-net variants. For one-step inference the pattern changes. Here, the FNO variants are on average more expensive than the U-nets. During rollout inference, however, the differences between models become smaller. This is likely because the fixed overhead of model loading constitutes a larger fraction of the total cost. These discrepancies highlight the importance of considering both training and inference costs rather than focusing on a single stage of the modelling pipeline. Although carbon footprint and runtime are naturally correlated, they are not strictly proportional across models (Fig. 6).
@@ 5 Discussions
Figure 6: Training runtime versus carbon footprint. The plot shows that energy consumption does not scale one-to-one with runtime.
Fig. 6 emphasises the importance of explicit energy or carbon measurement. Runtime alone is an unreliable proxy, as different hardware operations have distinct power profiles and a model that trains faster does not necessarily train cheaper [5].
@@
Our study demonstrates that physics inductive biases impact the trade-off between efficacy and efficiency. In the simple PINN experiment (Sec. 1), combining data with physics yielded the model dominating the Pareto front, achieving lower test errors at lower carbon cost compared to both the purely data-driven and purely physics-driven baselines.
@@
In the incompressible shear flow setting, we observe clear differences in how the physics priors affect both predictive performance and carbon footprint across training and inference. Incorporating weak physics-inductive biases in the form of periodic padding, improves performance but increases total training emission, while its effects on training carbon efficiency and one-step inference are model-dependent: increasing both in one case (UNet-CP) and decreasing both in the other (UNet-CN-CP). Models with stronger physics inductive biases achieve the lowest training footprints, but this comes at the expense of higher inference costs compared to U-net variants. UFNO achieves one of the most favourable overall profiles, combining strong predictive accuracy with a low training footprint. Its inference cost, however, remains nearly twice that of the best-performing U-net variant (UNet-CN-CP). The results underscore the importance of evaluating training and inference costs together. A model that trains cheaply may prove inefficient when repeatedly deployed, while others entail higher upfront cost but amortise more effectively in deployment. Amortising development costs across the model lifecycle is one way to factor in these considerations [33].
@@
Several key points follow from these results. First, model selection cannot be based on efficacy alone, as models with similar predictive accuracy can differ substantially in carbon emissions. Second, once both training and inference costs are accounted for, identifying a single “best” model becomes less straightforward. A model with a small carbon footprint during training may be unsuitable in deployment, making carbon-aware evaluation across the full lifecycle essential.
@@
The implications extend well beyond our test-bed. Many scientific and engineering problems are governed by well-understood structures, and overlooking them is wasteful. At the same time, tailoring models too closely to the specific physics can limit applicability and can be challenging for complex systems (e.g., PINNs). This motivates the development of general-purpose models such as FNOs, which allow physics priors to be incorporated more easily while retaining flexibility across tasks. However, alignment with the underlying physics does not guarantee carbon efficiency. Physics should not be incorporated at any cost, but only where it provides clear benefits in terms of predictive performance or carbon efficiency. While our work centres on physics-inductive biases, the same logic extends to other forms of prior knowledge: any regularising bias that constrains the learning task appropriately can potentially offer similar trade-offs between efficacy and efficiency.
@@
Ultimately, our results argue for a shift in mindset. Progress in ML should not be equated with ever-larger datasets, models, and compute budgets, but rather evaluated in terms of the value gained relative to the cost incurred [34]. Learning everything from data while ignoring domain knowledge reflects an abundance mindset, in which resources are treated as unlimited. A more sustainable alternative is to exploit what we already know by embedding this knowledge into models. This work offers a step towards understanding how prior knowledge shapes the efficacy-efficiency trade-offs, and argues for evaluating these explicitly.
@@
We did not perform full hyperparameter sweeps for the models. As detailed in App. 0.C, all configurations follow settings from prior work tuned on several PDE datasets, including the shear-flow dataset used here. This may influence the observed trade-offs. Furthermore, we restricted our study to spatio-temporal forecasting. While this limits the scope of the contribution, the central message about taking efficiency considerations into account by including appropriate inductive biases is applicable across ML tasks. Finally, the evaluation metrics focus on predictive accuracy rather than physical fidelity. Physics-based metrics would more directly assess whether the models capture underlying structures or conservation properties.
@@ 6 Conclusions
The primary message in this work, backed by empirical evidence, is to increase resource-awareness when designing ML experiments. While this can be accomplished in several ways [35], we focus on the role of physics inductive biases and illustrate the different decisions practitioners have to consider. Using spatio- temporal forecasting of incompressible shear flow, we show how physics-inductive biases affect the trade-off between predictive performance and carbon footprint.
This work advances a perspective grounded in resource-awareness and calls for a shift in mindset. As a community, we should discourage the pursuit of marginal improvements in performance that come from disproportionately higher resource costs. With our work we have made the case to encourage resource-awareness by utilising existing domain knowledge and/or other inductive biases.
{credits}
6.0.1 Acknowledgements
@@
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
@@
Appendix 0.A Simple Experiment with PINNs
Figure 7: Target and training data for the viscous Burgers’ equation. a: Target data. b: Supervised samples. c: PDE residual samples. d: Initial and boundary condition points.
0.A.1 Experimental Setup
In this simple experiment, we explore PINNs [36] which embed governing PDEs directly into the training objective, thereby regularising the optimisation landscape and guiding solutions towards physical consistency. By minimizing PDE residuals, PINNs can be trained with little or no supervised data.
We study the solution of the viscous Burgers’ equation with Dirichlet boundary conditions on a 2D spatio-temporal grid:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$ (2)
$u(x,0)=-\sin(\pi x),$ (3)
$u(1,t)=u(-1,t)=0,$ (4)
for $x\in[-1,1]$, $t\in[0,1]$, and with fixed viscosity ($\nu=0.025$). The domain is discretized with $256$ spatial points and $100$ time steps (Fig. 7).
We compare three models that share the same underlying architecture (four hidden layers, width 50, tanh activations) but differ in their use of physics through the loss function. The purely data-driven NN minimizes the mean-squared-error (MSE) between predictions $\hat{u}$ and ground truth $u$:
$\mathcal{L}_{\text{NN}}=\mathcal{L}_{\text{data}}=\frac{1}{N_{\text{data}}}\sum_{i=1}^{N_{\text{data}}}(u_{i}-\hat{u}_{i})^{2}.$ (5)
The unsupervised U-PINN is trained solely on physics constraints, minimizing PDE residuals alongside deviations from initial and boundary conditions:
@@ 0.A.1 Experimental Setup
$=\frac{1}{N_{\text{pde}}}\sum_{i=1}^{N_{\text{pde}}}\left(\frac{\partial\hat{u}_{i}}{\partial t}+\hat{u}_{i}\frac{\partial\hat{u}_{i}}{\partial x}-\nu\frac{\partial^{2}\hat{u}_{i}}{\partial x^{2}}\right)^{2}$
@@
The semi-supervised S-PINN combines both objectives, with data and physics terms weighted equally:
$\mathcal{L}_{\text{S-PINN}}=\mathcal{L}_{\text{data}}+\mathcal{L}_{\text{pde}}+\mathcal{L}_{\text{ic}}+\mathcal{L}_{\text{bc}}.$ (7)
All models were trained using Adam optimizer followed by L-BFGS optimizer [37] with learning rate $10^{-3}$. Dataset sizes range from $10^{3}$ to $5\times 10^{3}$ training points for PINNs and up to $8\times 10^{3}$ for the NN. Data and code are adapted from [38].
0.A.2 Results
Both PINNs achieve consistently lower test losses at lower carbon footprints compared to the supervised NN (Fig. 1-b). While the NN has the lowest carbon footprint for a fixed number of epochs and data points, achieving a given accuracy level is up to $10$ times more carbon-efficient with the PINNs. The S-PINN, which combines data and physics, dominates the Pareto front, suggesting that the right balance between data and physics yields the most favourable trade-off between predictive performance and carbon cost.
Appendix 0.B Incompressible Shear Flow Dataset
This experiment is based on data from The Well [26], consisting of 2D incompressible Navier–Stokes dynamics with an additional passive tracer: